### v1

# Identification of gapless phases by squaring a twist operator

Abstract
We propose a general necessary condition for a spin chain with SO(3) spin-rotation symmetry to be gapped. Specifically, we prove that the ground state(s) of an SO(3)-symmetric gapped spin chain must be spin singlet(s), and the expectation value of the square of a twist operator asymptotically approaches unity in the thermodynamic limit, where finite-size corrections are inversely proportional to the square root of the system size. This theorem provides (i) supporting evidence for various conjectured gapped phases, and (ii) a sufficient criterion for identifying gapless spin chains. We verify our theorem by numerical simulations for a variety of spin models and show that it offers a novel efficient way to identify gapless phases in spin chains with spin-rotation symmetry.
Introduction.—
Identifying and classifying various phases of quantum many-body systems is a subject of great interest in statistical and condensed matter physics [1]. Characteristic features of those quantum systems are observed in their low-energy sector, e.g., ground-state wave function(s) and various low-lying excitations, and roughly speaking, the quantum many-body systems are classified into gapped and gapless phases based on their distinct low-energy spectra [2, 3].
However, theoretical determination of the presence or absence of an excitation gap in many-body systems can be a notoriously difficult task because the “gapped” and “gapless” phases are concepts defined in the thermodynamic limit. A useful clue is provided by the famous Lieb-Schultz-Mattis (LSM) theorem [4] and its various extensions [5, 6, 7, 8], stating that one-dimensional quantum spin models with a half-integral spin per unit cell cannot be gapped with a unique ground state in the presence of SO(3) spin-rotation symmetry and lattice translation symmetry. Nevertheless, the LSM-type theorems have two shortcomings. (a) Since they are no-go type theorems, they cannot make definitive statements when a required microscopic condition is not satisfied by lattice spin models (e.g., spin-1 or integral spin chains). (b) Even for a lattice model satisfying the microscopic conditions required by the LSM-type theorems, they cannot distinguish gapless models from gapped models with degenerate ground states. To resolve these conceptual and substantial difficulties remains an open question.
In this Letter, we present a general, necessary condition for spin chains with spin-rotation symmetry to be gapped, and show that it can be used to identify gapless phases. Important ingredients for our discussion are the twist operator [9, 6, 10, 11, 4, 12]
$\hat{U}=\exp\left(\frac{2\pi i}{L}\sum_{j=1}^{L}j\hat{S}^{z}_{j}\right)$ (1)
and the ground-state expectation value of its square, $\langle\hat{U}^{2}\rangle$. We rigorously prove that $\langle\hat{U}^{2}\rangle=1+\mathscr{O}(1/\sqrt{L})\rightarrow 1$ for any ground state of gapped SO(3)-symmetric spin chains with extended translation symmetry (c.f. Theorem 5). We note that, unlike the LSM theorem, our theorem does not require stringent conditions, besides the SO(3) symmetry, on microscopic data per unit cell of spin chains, so that it is applicable to spins of any size $S$. Furthermore, it is numerically tractable since an arbitrarily chosen ground state can be used to evaluate $\langle\hat{U}^{2}\rangle$. We perform density matrix renormalization group (DMRG) simulations for several spin models [13, 14, 15], and present both evidence for (potentially) gapped phases and identification of gapless phases, which is ensured by reliable finite-size scaling analysis (with truncation errors below $10^{-7}$).
Symmetry, locality, and extended translation.—
To set the stage, we introduce various notations and concepts that are essential in our discussion of spin chains. A spin-chain Hamiltonian $\mathcal{H}$ is defined on a one-dimensional lattice of length $L$, where spin $\bm{S}_{j}=(\hat{S}^{x}_{j},\hat{S}^{y}_{j},\hat{S}^{z}_{j})$ acts on the $(2s_{j}+1)$-dimensional Hilbert space on site $j$, under periodic boundary conditions (PBCs), $\bm{S}_{j+L}=\bm{S}_{j}$. The Hamiltonian possesses $\text{U}(1)\rtimes\mathbb{Z}_{2}\subset\mathrm{SO}(3)$ spin-rotation symmetry; the full SO(3) symmetry will be imposed later (from Theorem 3). Here, U(1) and $\mathbb{Z}_{2}$ spin-rotation symmetries are generated by $\hat{S}^{z}_{\mathrm{tot}}$ and $R_{\pi}^{x}\equiv\exp(i\pi\hat{S}^{x}_{\mathrm{tot}})$, respectively, where $\hat{S}^{\alpha}_{\mathrm{tot}}=\sum_{j=1}^{L}\hat{S}^{\alpha}_{j}$ ($\alpha=x,y,z$).
The Hamiltonian is assumed to be local, i.e., it admits a decomposition $\mathcal{H}=\sum_{j=1}^{L}h_{j}$, where each $h_{j}$ acts as the identity operator on the spins at a distance larger than $d_{L}$ from site $j$, with $d_{L}$ being upper bounded by some model-dependent $d$ for $L\gg 1$. Moreover, the local decomposition can always be chosen to respect the on-site symmetry [12],
$[\hat{S}^{z},h_{j}]=[R^{x}_{\pi},h_{j}]=0.$ (2)
Furthermore, the Hamiltonian is assumed to be invariant under a translation $\hat{T}$
$\hat{T}^{-1}\bm{S}_{j}\hat{T}=\bm{S}_{j+r},$ (3)
where an integer $r$ is a model-dependent minimum period, so that we can increase the system size $L$ as $L=Nr$ with $N\rightarrow\infty$ to define the thermodynamic limit. We call the translation by the minimum period $r$ extended translation. There is a useful relationship between $\hat{U}$ and $\hat{T}$:
$\hat{T}^{-1}\hat{U}\hat{T}=\hat{U}\exp\!\left(2\pi i\sum_{j=1}^{r}\hat{S}^{z}_{j}\right)\exp\!\left(-\frac{2\pi i}{L}r\hat{S}^{z}_{\mathrm{tot}}\right),$ (4)
where $\exp\!\left(2\pi i\sum_{j=1}^{r}\hat{S}^{z}_{j}\right)\in\{\pm 1\}$ is simply a $\mathbb{Z}_{2}$ phase. One should note that the period $r$ can be any integer and model-dependent, and we do not need to impose translation symmetry with $r=1$ or any other microscopic constraint on each unit cell, unlike the LSM theorem.
Gapped and gapless Hamiltonians.— For the sake of rigor, we introduce the concept of “gapped” and “gapless” following Refs. [3, 2]. The system is said to be gapped in the limit of the system size $L\rightarrow\infty$, when the following condition is met: the ground-state degeneracy $g_{L}$ of the $L$-series of Hamiltonians $\mathcal{H}_{L}$ is upper bounded by a finite integer $g$, and the energy gap $\Delta_{L}$ between the $g_{L}$ ground state(s) and the first excited states of $\mathcal{H}_{L}$ is lower bounded by a finite positive number $\Delta$. Otherwise, the system is gapless.
As preparations for our main statements, we prove several Lemmas.
Lemma 1.—
[…]
Proof: Following the authors’ previous work [12], we calculate $\Delta E\equiv\langle\Psi|\mathcal{H}|\Psi\rangle-\langle\psi|\mathcal{H}|\psi\rangle$ and use the locality of $\mathcal{H}$ and the U(1) symmetry of $h_{j}$ to obtain
$\Delta E\!=\!\sum_{j=1}^{L}\left\langle\psi\bigg{|}\left[h_{j},\frac{2\pi i}{L}\!\!\sum_{|m-j|\leq d}(m-j)\hat{S}^{z}_{m}\right]\bigg{|}\psi\right\rangle\!+\!\mathscr{O}\!\left(\frac{1}{L}\right)\!.$
Since $h_{j}$ is $R_{\pi}^{x}$-symmetric and $|\psi\rangle$ is a $R_{\pi}^{x}$-eigenstate, we can replace $\hat{S}^{z}_{m}$ by $\left[\hat{S}_{m}^{z}+R_{\pi}^{x}\hat{S}_{m}^{z}{\left(R_{\pi}^{x}\right)}^{-1}\right]/2$, which vanishes. This completes the proof.
Lemma 2.— For any eigenstate $|\psi_{\text{gs}}\rangle$ in the ground-state sector of a spin-chain Hamiltonian $\mathcal{H}$ with U(1)$\rtimes\mathbb{Z}_{2}$ symmetry, the twisted state $\hat{U}|\psi_{\text{gs}}\rangle$ can be written as
$\hat{U}|\psi_{\text{gs}}\rangle=|\Psi_{\text{gs}}\rangle+|\xi\rangle,$ (5)
where $|\Psi_{\text{gs}}\rangle$ is a state in the ground-state sector with almost a unit norm and $|\xi\rangle$ is an excited state with nearly a zero norm, i.e.,
$\langle\Psi_{\text{gs}}|\Psi_{\text{gs}}\rangle=1+\mathscr{O}\!\left(\frac{1}{L}\right)\!,\,\,\,\langle\xi|\xi\rangle=\mathscr{O}\!\left(\frac{1}{L}\right)\!,\,\,\,\langle\xi|\Psi_{\text{gs}}\rangle=0.$
Proof: We take an orthonormal basis $\{|\text{gs}_{j}\rangle|{j=1,2,\cdots,g_{L}}\}$ of the ground-state sector, where $|\text{gs}_{j}\rangle$ are simultaneously $R_{\pi}^{x}$-eigenstates. We write $|\psi_{\text{gs}}\rangle$ as
$|\psi_{\text{gs}}\rangle=\sum_{j=1}^{g}\gamma_{j}|\text{gs}_{j}\rangle,\quad\sum_{j=1}^{g}|\gamma_{j}|^{2}=1.$ (6)
For each $|\text{gs}\rangle_{j}$, Lemma 1 implies $\hat{U}|\text{gs}_{j}\rangle=|\Psi_{\text{gs$j$}}\rangle+|\xi_{j}\rangle$. Here it is important to note that the existence of a nonzero gap $\Delta$ ensures that the ground-state component $|\Psi_{\text{gs$j$}}\rangle$ has a norm $\langle\Psi_{\text{gs$j$}}|\Psi_{\text{gs$j$}}\rangle=1+\mathscr{O}(1/L)$ while $|\xi_{j}\rangle$ above the ground state(s) has a vanishing norm $\langle\xi_{j}|\xi_{j}\rangle=\mathscr{O}(1/L)$. Thus
$\hat{U}|\psi_{\text{gs}}\rangle=\sum_{j=1}^{g}\gamma_{j}\left(|\Psi_{\text{gs}j}\rangle+|\xi\rangle_{j}\right)\equiv|\Psi_{\text{gs}}\rangle+|\xi\rangle,$ (7)
where the first term on the right-hand side gives a state in the ground-state sector, and the norm of second term is given by $\langle\xi|\xi\rangle=\sum_{j,k=1}^{g}\gamma_{j}^{*}\gamma_{k}\langle\xi_{j}|\xi_{k}\rangle=\mathscr{O}(1/L)$. Consequently, the unitarity of $\hat{U}$ gives $\langle\Psi_{\text{gs}}|\Psi_{\text{gs}}\rangle=1+\mathscr{O}(1/L)$ completing the proof.
Theorem 3: Singlet criterion.— If spin-chain Hamiltonian $\mathcal{H}$ with extended translation symmetry and SO(3) spin-rotation symmetry is gapped, then any eigenstate in its ground-state sector is a spin singlet state $\bm{\hat{S}}_{\mathrm{tot}}=0$.
[…]
$\hat{T}|\text{gs}\rangle=\exp(iP_{0})|\text{gs}\rangle,$ (8)
where $P_{0}$ is its lattice momentum. Equation (4) gives
$\hat{T}\hat{U}|\text{gs}\rangle=\exp[i(P_{0}+\Delta P)]\hat{U}|\text{gs}\rangle,$ (9)
where $\Delta P$ takes a definite value in the following set:
$\Delta P\in\left\{-\frac{\pi r}{L},-\frac{2\pi r}{L},\pi-\frac{\pi r}{L},\pi-\frac{2\pi r}{L}\right\}.$ (10)
Thus, the state $\hat{U}|\text{gs}\rangle$ has the lattice momentum $P_{0}+\Delta P$ and is a normalized state in the ground-state sector in the thermodynamic limit, according to Lemma 2. Repeating this procedure generates the states $\hat{U}^{q}|\text{gs}\rangle$ for $q=1,2,3,\cdots$, a sequence of ground states with distinct momenta $P_{0}+q\Delta P$ in the thermodynamic limit. As a result, the number of these states in the ground-state sector is at least of the order of $\mathscr{O}(L/r)$, regardless of the specific value $\Delta P$ taken in Eq. (10). Since $r$ remains finite in the thermodynamic limit, the ground-state degeneracy must diverge $\propto L$, which contradicts the assumption of upper-bounded ground-state degeneracy in the definition of gapped Hamiltonians. This completes the proof of Theorem 3.
Theorem 3 provides a powerful criterion for determining gaplessness in SO(3)-symmetric spin chains with extended translation symmetry. For instance, it directly ensures the gaplessness of half-integer spin chains of odd lengths, since such systems do not admit singlet states in its Hilbert space.
Corollary 4.—
Any SO(3)-symmetric spin chain with half-integral spin on each site with an extended periodicity must be gapless if the thermodynamic limit is taken as $L=2N+1$ with $N\to\infty$.
The consideration that the entire odd-$L$ half-integer chain, as a zero-dimensional quantum mechanical system, forms a projective representation of SO(3) symmetry only implies that the ground-state energy (actually any energy) must be at least doubly degenerate. The gaplessness imposed by Corollary 4 is genuinely a many-body effect beyond such a zero-dimensional single-body viewpoint.
With these preparatory Corollaries and Theorems, we are now ready to propose and prove our main theorem.
Theorem 5: $\langle\hat{U}^{2}\rangle$-criterion.—
[…]
$\langle\text{GS}|\hat{U}^{2}|\text{GS}\rangle=1+\mathscr{O}(1/\sqrt{L})\rightarrow 1.$ (11)
Proof: Lemma 2 implies that
$\hat{U}|\text{GS}\rangle=|\Psi_{\text{GS}}\rangle+|\gamma\rangle,$ (12)
with $\langle\Psi_{\text{GS}}|\Psi_{\text{GS}}\rangle=1+\mathscr{O}(1/L)$ and $\langle\gamma|\gamma\rangle=\mathscr{O}(1/L)$. It follows from the identity $\hat{U}R_{\pi}^{x}\hat{U}(R_{\pi}^{x})^{-1}=1$ that
$1=\langle\text{GS}|\hat{U}R_{\pi}^{x}\hat{U}(R_{\pi}^{x})^{-1}|\text{GS}\rangle.$ (13)
Since any state in the ground-state sector is an SU(2) singlet by Theorem 3, $R_{\pi}^{x}|\text{GS}\rangle=(R_{\pi}^{x})^{-1}|\text{GS}\rangle=|\text{GS}\rangle$ and $R_{\pi}^{x}|\Psi_{\text{GS}}\rangle=(R_{\pi}^{x})^{-1}|\Psi_{\text{GS}}\rangle=|\Psi_{\text{GS}}\rangle$. Thus, Eq. (13) becomes
11 11 $=$ $\langle\text{GS}|\hat{U}R_{\pi}^{x}(|\Psi_{\text{GS}}\rangle+|\gamma\rangle)$ (14)
$=$ $\langle\text{GS}|\hat{U}|\Psi_{\text{GS}}\rangle+\langle\text{GS}|\hat{U}R_{\pi}^{x}|\gamma\rangle,$
where the second term is of order $\mathscr{O}(1/\sqrt{L})$ since $\sqrt{\langle\gamma|\gamma\rangle}=\mathscr{O}(1/\sqrt{L})$ and the unitary operator $\hat{U}R_{\pi}^{x}$ is norm preserving. On the other hand,
$\langle\text{GS}|\hat{U}^{2}|\text{GS}\rangle$ $=$ $\langle\text{GS}|\hat{U}(|\Psi_{\text{GS}}\rangle+|\gamma\rangle)$ (15)
$=$ $\langle\text{GS}|\hat{U}|\Psi_{\text{GS}}\rangle+\langle\text{GS}|\hat{U}|\gamma\rangle,$
where the second term is $\mathscr{O}(1/\sqrt{L})$. Combing Eqs. (14) and (15), we obtain
$\langle\text{GS}|\hat{U}^{2}|\text{GS}\rangle$ $=$ $1-\langle\text{GS}|\hat{U}R_{\pi}^{x}|\gamma\rangle+\langle\text{GS}|\hat{U}|\gamma\rangle$ (16)
$=$ $1+\mathscr{O}(1/\sqrt{L}),$
which completes the proof of Theorem 5.
Theorem 5 can be formally generalized by defining
$\hat{U}_{\bm{n}}\equiv\exp\left[i\frac{2\pi}{L}\sum_{j=1}^{L}j\left(\bm{n}\cdot\bm{S}_{j}\right)\right],$ (17)
with a unit 3-vector $\bm{n}$  [^1] One should note that $\hat{U}_{\bm{n}+\bm{m}}\neq\hat{U}_{\bm{n}}\hat{U}_{\bm{m}}$ in general.. For any $\bm{n}$, there exists an SU(2) spin-rotation transformation $R_{\bm{n}}$ such that $\hat{S}^{z}_{j}=R_{\bm{n}}\left(\bm{n}\cdot\bm{S}_{j}\right)R_{\bm{n}}^{-1}$ or $\hat{U}=R_{\bm{n}}\hat{U}_{\bm{n}}R_{\bm{n}}^{-1}$. Noticing $R_{\bm{n}}|\text{GS}\rangle=|\text{GS}\rangle$ and unitarity of $\hat{U}_{\bm{n}}$ and $R_{\bm{n}}$, we obtain the following formal extension.
Corollary 6.—Under the same condition of Theorem 5,
$\langle\text{GS}|\hat{U}_{\bm{n}_{1}}\hat{U}_{\bm{n}_{2}}\cdots\hat{U}_{\bm{n}_{p}}|\text{GS}\rangle=1+\mathscr{O}(1/\sqrt{L}),$ (18)
for any even $p\in 2\mathbb{Z}$ and unit 3-vectors $\bm{n}_{1,2,\cdots,p}$.
Identifying gaplessness and evidence of gappedness.—
Since Theorem 5 provides a necessary quantitative condition for an SO(3)-symmetric Hamiltonian to be gapped, it can be utilized to identify gapless Hamiltonian as well as provide supporting evidence for various potentially gapped Hamiltonian.
Figure 1: Results for $\langle\hat{U}^{2}\rangle$ in the Majumdar-Ghosh model. (a) For even chain lengths, DMRG data correspond to the first ground state at each system size $L$. The expectation values scale linearly versus $1/L$ [see Eq. (20)], with intercept of 0.996(1) in the thermodynamic limit. The analytical results for two exact ground states are also showed into pink and blue lines (see Supplementary Materials [17]). (b) For odd chain lengths, the expectation value remains below $10^{-4}$ for all system sizes considered.
Let us begin with the $S=1/2$ Majumdar-Ghosh (M-G) model [18, 19, 20]
$\mathcal{H}_{\text{M-G}}=\sum_{j=1}^{L}\bm{S}_{j}\cdot\bm{S}_{j+1}+\frac{1}{2}\sum_{j=1}^{L}\bm{S}_{j}\cdot\bm{S}_{j+2},$ (19)
with length $L$ and PBCs. For even chain lengths $L\in 2\mathbb{Z}$, its lowest-energy sector is spanned by $|\psi_{\text{M-G}}\rangle=|(1,2)(3,4)\ldots(L-1,L)\rangle$ and $|\phi_{\text{M-G}}\rangle=|(2,3)(4,5)\ldots(L,1)\rangle$, where $(m,n)$ denotes a singlet formed by the spins on the sites $m$ and $n$ [21, 22, 23, 24]. Each singlet contributes a factor $\cos[(2\pi/L)(m-n)]$ in evaluating $\langle\text{GS}|\hat{U}^{2}|\text{GS}\rangle$, yielding [17]
$\langle\Psi_{\text{M-G}}^{\text{(even)}}|\hat{U}^{2}|\Psi_{\text{M-G}}^{\text{(even)}}\rangle=1-\frac{\pi^{2}}{L}+\mathscr{O}\left(\frac{1}{L^{2}}\right).$ (20)
for any lowest-energy state $|\Psi_{\text{M-G}}^{\text{(even)}}\rangle$. Although this result is exact, we also show DMRG results for system sizes up to $L=192$ [Fig 1 (a)], which yields an intercept of 0.996(1) by extrapolating to the thermodynamic limit.
By contrast, for odd chain lengths $L\in 2\mathbb{Z}+1$ [23], there is no available exact ground-state wavefunction. The DMRG simulations reveal that $\langle\Psi_{\text{M-G}}^{\text{(odd)}}|\hat{U}^{2}|\Psi_{\text{M-G}}^{\text{(odd)}}\rangle$ remains below $10^{-4}$ for all systems sizes under consideration, indicating the gaplessness. It is in accordance with the conclusion from Corollary 4 that odd-$L$ MG chains are gapless.
We also apply the $\langle\hat{U}^{2}\rangle$-criterion to spin-$S$ Heisenberg antiferromagnetic (HAF) chains
$\mathcal{H}_{\text{HAF}}=\sum_{j}\bm{S}_{j}\cdot\bm{S}_{j+1}.$ (21)
The $S=\frac{1}{2}$ HAF chain is known to be gapless from the Bethe ansatz solutions [25, 26], while the $S=1$ HAF chain is believed to be gapped [27, 28, 29], as supported by numerical evidence [30, 31, 32]. We calculate $\langle\hat{U}^{2}\rangle$ and fit the DMRG data, up to system size $L=128$ for both $S=\frac{1}{2}$ and $1$, to the form
$\langle\text{GS}|\hat{U}^{2}|\text{GS}\rangle=a-\frac{b}{L^{k}}.$ (22)
We obtain $a=-0.053(5)$, $b=0.284(5)$, and $k=0.481(4)$ for $S=\frac{1}{2}$ in Fig. 2 (a), which signals gaplessness. For $S=1$, we find the intercepts $a=1.003(9)$, with $b=11.568(4)$, $k=0.853(0)$ as shown in Fig. 2 (b), which is consistent with the Haldane conjecture [27, 28].
Figure 2: Results for $\langle\hat{U}^{2}\rangle$ as a function of $1/L$ for spin-$S=1/2$ (left) and spin-$S=1$ HAF chains (right). Each dataset is fit by Eq. (22). For $S=1/2$, the fitting yields $b=0.284(5)$, $k=0.481(4)$, and $a=-0.053(5)$; for $S=1$, the fitting gives $b=11.568(4)$, $k=0.853(0)$, and $a=1.003(9)$.
We also consider the spin-1 Affleck-Kennedy-Lieb-Tasaki (AKLT) model [33]
$\mathcal{H}_{\text{AKLT}}=\sum_{j=1}^{L}\bm{S}_{j}\cdot\bm{S}_{j+1}+\frac{1}{3}\sum_{j=1}^{L}\left(\bm{S}_{j}\cdot\bm{S}_{j+1}\right)^{2}.$ (23)
Our DMRG results confirm its gapped behavior [33, 23]; the extrapolation to the $L\to\infty$ limit with the least-square fitting yields the intercept $a=1.000(3)$, with $b=8.073(2)$, $k=0.965(2)$ as shown in Fig. 3.
Figure 3: Results for $\langle\hat{U}^{2}\rangle$ as a function of $1/L$ for the AKLT chain. The DMRG results are fitted to Eq. (22), where the fitting yields $a=1.000(3)$ , $b=8.073(2)$, and $k=0.965(2)$.
Conclusions and discussions—
In this work, we present a rigorous criterion for gaplessness of spin-rotation symmetric spin chains by proving that, for any gapped Hamiltonian, (a) its ground state(s) must be SU(2)-singlet; (b) the ground-state expectation value of $\hat{U}^{2}$ converges to unity in the thermodynamic limit with the finite-size corrections upper bounded by $\mathscr{O}(1/\sqrt{L})$. We apply our theorem to provide supporting evidence for potentially gapped spin models and identify various gapless spin chains by numerics. Unfortunately, our theorem assumes the full SO(3) spin-rotation symmetry [34], which cannot be applied to, e.g., XXZ models [35, 36]. The extension of the current result to spin models with discrete spin-rotation symmetry is left for future work.
Acknowledgements.—
The calculations of this works are performed using the ITensor Julia library (version 0.7) [15]. The computations in this Letter were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The authors thank Linhao Li and Masaki Oshikawa for helpful discussions. The work of Y. Y. was supported by the National Key Research and Development Program of China (Grant No. 2024YFA1408303), the National Natural Science Foundation of China (Grant No. 12474157), the sponsorship from Yangyang Development Fund, and Xiaomi Young Scholars Program. The work of A.F. was supported in part by JST CREST (Grant No. JPMJCR19T2).
Appendix A Supplemental Materials
A.1 Analytical results of $\langle\Psi_{\text{M-G}}|\hat{U}^{2}|\Psi_{\text{M-G}}\rangle$ for Majumdar-Ghosh model defined on even-length chains
As stated in the text, there are two degenerated VBS ground states when the length is chosen as an even number $L\in 2\mathbb{Z}$,
$|\Psi_{\text{M-G}}^{(1)}\rangle$ $=$ $|(1,2)(3,4)\ldots(L-1,L)\rangle,$ (24)
$|\Psi_{\text{M-G}}^{(2)}\rangle$ $=$ $|(2,3)(4,5)\ldots(L,1)\rangle,$ (25)
where $(m,n)$ denotes singlet localized on sites $m$ and $n$,
$|(m,n)\rangle=\frac{1}{\sqrt{2}}\left(|\uparrow_{m}\downarrow_{n}\rangle-|\downarrow_{m}\uparrow_{n}\rangle\right).$ (26)
One should note that
$\langle\Psi_{\text{M-G}}^{(2)}|\Psi_{\text{M-G}}^{(1)}\rangle\propto\left(\frac{1}{\sqrt{2}}\right)^{L/2}=\mathscr{O}\left(\frac{1}{L^{+\infty}}\right),$ (27)
which enables us to consider that they are orthogonal in the following $L$-scaling.
Since each singlet localized on two different sites is the eigenstate of $\hat{S}^{z}$ on those sites, one can consider the contribution from each singlet $(m,n)$ more generally, for expectation value of $\hat{U}^{q}$,
$\langle(m,n)|\hat{U}^{q}|(m,n)\rangle$ $=$ $\frac{1}{2}\left\{\exp\left[i\frac{2\pi q}{L}\left(\frac{m}{2}-\frac{n}{2}\right)\right]+\exp\left[i\frac{2\pi q}{L}\left(\frac{n}{2}-\frac{m}{2}\right)\right]\right\}$ (28)
$=$ $\cos\!\left(q\pi\frac{m-n}{L}\right),$
which for our case $q=2$, equals to
$\langle(m,n)|\hat{U}^{2}|(m,n)\rangle=\cos\left(2\pi\frac{m-n}{L}\right).$ (29)
Thus we can calculate out for the VBS states
$\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle$ $=$ $\left[\cos\left(\frac{2\pi}{L}\right)\right]^{L/2},$ (30)
$\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle$ $=$ $\left[\cos\left(\frac{2\pi}{L}\right)\right]^{L/2-1}\cos\left[\frac{2\pi}{L}(L-1)\right].$ (31)
By taking thermodynamic limit $L\rightarrow\infty$, we obtain
$\lim_{L\rightarrow\infty}\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle$ $=$ $1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}-\frac{4\pi^{4}+\pi^{6}}{6L^{3}}+\mathscr{O}\left(\frac{1}{L^{4}}\right),$ (32)
$\lim_{L\rightarrow\infty}\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle$ $=$ $1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}-\frac{(-2/3)\pi^{4}+(1/6)\pi^{6}}{6L^{3}}+\mathscr{O}\left(\frac{1}{L^{4}}\right).$ (33)
Additionally,
$\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle\propto\left(\frac{1}{\sqrt{2}}\right)^{L/2}=\mathscr{O}\left(\frac{1}{L^{+\infty}}\right).$ (34)
We observe that the coefficients of $1/L$ and $1/L^{2}$ terms are the same for these two VBS ground states, resulting in the same coefficients of leading term for an arbitrary normalized ground state $|\Psi_{\text{M-G}}\rangle\equiv\alpha|\Psi_{\text{M-G}}^{(1)}\rangle+\beta|\Psi_{\text{M-G}}^{(2)}\rangle$ spanned by them, with $|\alpha|^{2}+|\beta|^{2}=1$,
$\lim_{L\rightarrow\infty}\langle\Psi_{\text{M-G}}|\hat{U}^{2}|\Psi_{\text{M-G}}\rangle$ $=$ $\lim_{L\rightarrow\infty}\left(|\alpha|^{2}\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle+|\beta|^{2}\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle\right)$ (35)
$=$ $\left(|\alpha|^{2}+|\beta|^{2}\right)\left[1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}+\mathscr{O}\left(\frac{1}{L^{3}}\right)\right]$
$=$ $1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}+\mathscr{O}\left(\frac{1}{L^{3}}\right).$